### v1

# PolyKAN: A Polyhedral Analysis Framework for Provable and Minimal KAN Compression

Abstract
Kolmogorov-Arnold Networks (KANs) have emerged as a promising alternative to traditional Multi-Layer Perceptrons (MLPs), offering enhanced interpretability and a strong mathematical foundation. However, their parameter efficiency remains a significant challenge for practical deployment. This paper introduces PolyKAN, a novel theoretical framework for KAN compression that provides formal guarantees on both model size reduction and approximation error. By leveraging the inherent piecewise polynomial structure of KANs, we formulate the compression problem as one of optimal polyhedral region merging. We establish a rigorous polyhedral characterization of KANs, develop a complete theory of $\epsilon$-equivalent compression, and design an optimal dynamic programming algorithm that guarantees minimal compression under specified error bounds. Our theoretical analysis demonstrates that PolyKAN achieves provably minimal compression while maintaining strict error control, with polynomial-time complexity in all network parameters. The framework provides the first formal foundation for KAN compression with mathematical guarantees, opening new directions for efficient deployment of interpretable neural architectures.
Keywords: Kolmogorov-Arnold Networks, Model Compression, Polyhedral Analysis, Provable Optimization, Approximation Theory, Computational Complexity
1 Introduction
The pursuit of interpretable yet powerful neural network architectures has led to renewed interest in mathematically grounded alternatives to traditional black-box models. Kolmogorov-Arnold Networks (KANs) [1], inspired by the celebrated Kolmogorov-Arnold representation theorem [2], have recently emerged as a compelling framework that replaces fixed activation functions with learnable spline transformations. While KANs demonstrate remarkable interpretability and empirical performance, their parameter efficiency presents a substantial obstacle to practical adoption, as each network connection requires an independent spline function with multiple parameters.
Existing approaches to neural network compression—including pruning [3], knowledge distillation [4], and architecture search—largely rely on heuristic strategies without formal guarantees. For KANs specifically, the absence of theoretically grounded compression methods represents a significant gap in the literature. The spline-based foundation of KANs, however, provides a unique opportunity for rigorous analysis through polyhedral theory, as each spline naturally partitions its input domain into piecewise polynomial regions.
In this paper, we present PolyKAN, a comprehensive theoretical framework for KAN compression with provable guarantees. Our work is distinguished by three fundamental contributions:
  1. 1.
    We establish a complete polyhedral characterization of KANs, demonstrating that their input space partitions exhibit a highly regular axis-aligned structure—in stark contrast to the arbitrary polyhedral complexes of ReLU networks. This structural regularity enables rigorous theoretical analysis.
  2. 2.
    We develop a formal theory of $\epsilon$-equivalent compression for KANs, providing necessary and sufficient conditions for region merging while maintaining strict error bounds. Our analysis includes precise error propagation bounds through multiple network layers.
  3. 3.
    We design and analyze an optimal dynamic programming algorithm that guarantees minimal compression under specified error constraints. We prove polynomial-time complexity and establish formal guarantees for both correctness and optimality.
The PolyKAN framework transforms KAN compression from a heuristic process to a mathematically rigorous optimization problem with verifiable guarantees. By bridging spline theory, polyhedral geometry, and algorithm design, we provide a foundation for trustworthy compression of interpretable neural architectures.
2 Background and Related Work
2.1 Kolmogorov-Arnold Networks and Spline Theory
The Kolmogorov-Arnold representation theorem [2] states that any multivariate continuous function can be represented as a finite composition of univariate functions and additions. KANs [1] operationalize this theorem by placing learnable univariate functions (typically B-splines) on network edges while limiting nodes to summation operations. This architectural shift from traditional MLPs provides both mathematical interpretability and empirical advantages in certain function approximation tasks.
Formally, a KAN layer transforms its input $\mathbf{x}\in\mathbb{R}^{n}$ to output $\mathbf{y}\in\mathbb{R}^{m}$ through the relation:
$y_{j}=\sum_{i=1}^{n}s_{j,i}(x_{i}),\quad j=1,\ldots,m$ (1)
where each $s_{j,i}:\mathbb{R}\to\mathbb{R}$ is a spline function parameterized by knot positions and polynomial coefficients.
The approximation properties of splines are well-understood [5], with deep connections to function space theory. Recent work [1] has begun exploring the implications of these properties for neural network design, but a comprehensive theoretical framework for KAN optimization remains underdeveloped.
2.2 Neural Network Compression
Model compression techniques aim to reduce the computational and memory requirements of neural networks without significant performance degradation. Pruning methods [3, 10] remove parameters or connections based on various importance criteria, while knowledge distillation [4] trains compact student networks to mimic larger teacher models. Neural architecture search [11] automates the design of efficient network structures.
Despite empirical successes, these approaches largely lack formal guarantees. Theoretical work on network compression has typically focused on simplified settings [12] or provided only asymptotic guarantees. The lottery ticket hypothesis [10] offers intriguing insights but falls short of providing constructive compression algorithms with bounded error.
2.3 Polyhedral Theory in Deep Learning
Polyhedral theory provides powerful tools for analyzing piecewise linear neural networks. A substantial body of work has studied the linear regions of ReLU networks [6, 7, 8], establishing connections between region counts and network expressivity. Mixed-integer programming formulations [9] have been developed for verifying properties of ReLU networks.
However, the polyhedral structure of KANs differs fundamentally from that of ReLU networks. While ReLU networks partition their input space with arbitrarily oriented hyperplanes, KAN partitions are axis-aligned due to their spline-based construction. This structural regularity makes KANs particularly amenable to polyhedral analysis and enables the development of compression algorithms with strong theoretical guarantees.
3 Polyhedral Characterization of KANs
3.1 Univariate Spline Partitions
We begin by formalizing the polyhedral structure of individual spline functions, which form the building blocks of KANs.
Definition 1 (Spline Polyhedron).
For a B-spline function $s:[a,b]\to\mathbb{R}$ with knot sequence $t_{0},t_{1},\ldots,t_{k}$ where $a=t_{0}<t_{1}<\cdots<t_{k}=b$, the spline polyhedron is a tuple $\mathcal{P}_{s}=(R,P)$, where:
  • $R=\{R_{1},R_{2},\ldots,R_{k}\}$ is a set of regions with $R_{i}=\{x\in\mathbb{R}:t_{i-1}\leq x\leq t_{i}\}$
  • $P=\{p_{1},p_{2},\ldots,p_{k}\}$ is a set of polynomial functions, with $p_{i}$ defined on $R_{i}$
This definition captures the piecewise polynomial nature of spline functions, where each region corresponds to an interval between consecutive knots.
3.2 Multi-Layer KAN Partitions
For multi-layer KANs, the polyhedral structure emerges from the composition of spline functions across layers. Consider an $L$-layer KAN with architecture $n_{0},n_{1},\ldots,n_{L}]$, where $n_{l}$ denotes the number of nodes in layer $l$.
Theorem 1 (KAN Polyhedral Region Structure).
The input space partition of an $L$-layer KAN is the refinement of partitions induced by spline functions across all layers. The total number of linear regions satisfies:
$N_{\text{regions}}\leq\prod_{l=1}^{L}\prod_{i=1}^{n_{l-1}}\prod_{j=1}^{n_{l}}(k^{(l)}_{j,i}-1)$ (2)
where $k^{(l)}_{j,i}$ is the number of knots in spline function $s^{(l)}_{j,i}$ connecting node $i$ in layer l1l-1 to node $j$ in layer $l$.
Proof.
Each spline function $s^{(l)}_{j,i}$ partitions its one-dimensional input space into at most $(k^{(l)}_{j,i}-1)$ intervals. Since these partitions act independently along different dimensions at each layer, the Cartesian product of these partitions yields the region structure for that layer’s output space. The composition across layers refines these partitions, leading to the product bound. ∎
3.3 Structural Properties of KAN Polyhedra
The polyhedral structure of KANs exhibits three fundamental properties that distinguish them from ReLU networks and enable our compression theory.
Lemma 1 (Axis-Alignment).
All boundary hyperplanes of KAN polyhedral regions are axis-aligned. That is, each boundary can be expressed as $x_{d}=c$ for some coordinate $d$ and constant $c$.
Proof.
The knot points of each spline function $s^{(l)}_{j,i}(x_{i})$ define partitions of the form $x_{i}=t_{m}$, which are hyperplanes orthogonal to the $i$-th coordinate axis. The composition of such axis-aligned partitions preserves axis-alignment. ∎
Lemma 2 (Rectangular Structure).
Every KAN polyhedral region is a axis-aligned rectangle (the Cartesian product of intervals).
Proof.
Following from the axis-alignment property, the intersection of half-spaces defined by axis-aligned hyperplanes necessarily produces rectangular regions. ∎
Lemma 3 (Function Regularity).
Within each polyhedral region, a KAN is a smooth multivariate polynomial function. Across region boundaries, the KAN maintains continuity (for linear B-splines) or higher-order smoothness (for higher-degree B-splines).
These structural properties significantly simplify polyhedral analysis compared to ReLU networks, where regions can be arbitrary convex polyhedra and boundaries have arbitrary orientations.
4 Provable KAN Compression Theory
4.1 Formal Problem Statement
We now formalize the KAN compression problem with provable guarantees.
Definition 2 ($\epsilon$-Equivalent Compression).
Given a KAN network $\mathcal{N}:\mathcal{X}\to\mathbb{R}$ and error tolerance $\epsilon>0$, a compressed network $\mathcal{N}^{\prime}$ is an $\epsilon$-equivalent compression of $\mathcal{N}$ if:
$\|\mathcal{N}-\mathcal{N}^{\prime}\|_{\infty}=\sup_{x\in\mathcal{X}}|\mathcal{N}(x)-\mathcal{N}^{\prime}(x)|\leq\epsilon$ (3)
Our objective is to find an $\epsilon$-equivalent compression $\mathcal{N}^{\prime}$ with minimal number of parameters (or equivalently, minimal total knots across all spline functions).
4.2 Region Mergability Theory
The foundation of our compression approach is the theory of merging adjacent polyhedral regions while controlling approximation error.
Definition 3 (Region Mergability).
Let $R_{i}$ and $R_{j}$ be two adjacent polyhedral regions in a KAN, with corresponding polynomial functions $p_{i}$ and $p_{j}$. The regions are $\epsilon$-mergable if there exists a single polynomial $p_{ij}$ such that:
$\max\left\{\max_{x\in R_{i}}|p_{i}(x)-p_{ij}(x)|,\max_{x\in R_{j}}|p_{j}(x)-p_{ij}(x)|\right\}\leq\epsilon$ (4)
This definition captures the intuition that we can replace two different polynomials defined on adjacent regions with a single polynomial that approximates both within tolerance $\epsilon$.
Theorem 2 (Knot Elimination Condition).
Let $t_{m}$ be an interior knot of spline function $s(x)$, with adjacent regions $R_{m-1}=[t_{m-1},t_{m}]$ and $R_{m}=[t_{m},t_{m+1}]$. If $R_{m-1}$ and $R_{m}$ are $\epsilon$-mergable, then knot $t_{m}$ can be eliminated while maintaining $\epsilon$-equivalence.
Proof.
If $R_{m-1}$ and $R_{m}$ are $\epsilon$-mergable, there exists a polynomial $p$ such that $|s(x)-p(x)|\leq\epsilon$ for all $x\in R_{m-1}\cup R_{m}$. By replacing the original piecewise representation with $p$ on the merged region $t_{m-1},t_{m+1}]$, we obtain a new spline function $\tilde{s}$ that satisfies $\|s-\tilde{s}\|_{\infty}\leq\epsilon$ while eliminating one knot. ∎
4.3 Error Propagation Analysis
For multi-layer KAN compression, we must understand how compression errors propagate through the network.
Lemma 4 (Single-Layer Error Propagation).
Consider a KAN layer $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$ with:
$y_{j}=\sum_{i=1}^{n}s_{j,i}(x_{i}),\quad j=1,\ldots,m$
If each spline $s_{j,i}$ is compressed to $\tilde{s}_{j,i}$ with $\|s_{j,i}-\tilde{s}_{j,i}\|_{\infty}\leq\delta$, then:
$\|\Phi-\tilde{\Phi}\|_{\infty}\leq n\cdot\delta$
Proof.
For any input $\mathbf{x}$ and output dimension $j$:
$|\Phi_{j}(\mathbf{x})-\tilde{\Phi}_{j}(\mathbf{x})|$ $=\left|\sum_{i=1}^{n}\left(s_{j,i}(x_{i})-\tilde{s}_{j,i}(x_{i})\right)\right|$
$\leq\sum_{i=1}^{n}|s_{j,i}(x_{i})-\tilde{s}_{j,i}(x_{i})|$
$\leq\sum_{i=1}^{n}\delta=n\delta$
Taking the supremum over all $\mathbf{x}$ and maximum over $j$ completes the proof. ∎
This lemma enables hierarchical error budgeting across network layers. Given a global error tolerance $\epsilon$, we can allocate error budgets $\delta_{l}$ to each layer $l$ such that the cumulative effect satisfies the global bound.
4.4 Optimal Compression Formulation
We now formulate KAN compression as a combinatorial optimization problem. Let $z_{l,i,j,m}\in\{0,1\}$ indicate whether the $m$-th knot of spline $s^{(l)}_{j,i}$ is retained, and let $\delta_{l,i,j}$ be the error budget allocated to this spline.
The optimal compression problem becomes:
$\min$ $\sum_{l=1}^{L}\sum_{i=1}^{n_{l-1}}\sum_{j=1}^{n_{l}}\sum_{m=1}^{k_{l,i,j}}z_{l,i,j,m}$ (5)
s.t. $\epsilon$-equivalence constraints
Knot sequence feasibility
$z_{l,i,j,m}\in\{0,1\}$
While this integer programming formulation captures the optimality structure, its direct solution is computationally challenging. In the next section, we develop an efficient dynamic programming algorithm that exploits the problem structure to achieve optimality with polynomial complexity.
5 The PolyKAN Compression Algorithm
5.1 Algorithm Overview
We present the PolyKAN compression algorithm, which systematically reduces KAN complexity while guaranteeing $\epsilon$-equivalence. The algorithm operates layer by layer, compressing each spline function optimally using dynamic programming.
Algorithm 1 PolyKAN Provable Compression Algorithm
1:Original KAN network $\mathcal{N}$, global error bound $\epsilon$
2:Compressed KAN network $\mathcal{N}^{\prime}$ with $\|\mathcal{N}-\mathcal{N}^{\prime}\|_{\infty}\leq\epsilon$
[…]
4:for $l\leftarrow 1$ to $L$ do
5:  for each spline function $s$ in layer $l$ do
6:   $\tilde{s}\leftarrow\text{OptimalSplineCompression}(s,\delta_{l})$ $\triangleright$ Algorithm 2
7:  end for
8:end for
[…]
10:return $\mathcal{N}^{\prime}$
The error allocation in Step 1 distributes the global error budget $\epsilon$ across layers according to the error propagation analysis, ensuring that the cumulative effect satisfies the global bound.
5.2 Optimal Spline Compression
The core of PolyKAN is the optimal compression of individual spline functions via dynamic programming.
Algorithm 2 OptimalSplineCompression
1:function OptimalSplineCompression($s$, $\delta$)
2:  $T\leftarrow[t_{0},t_{1},\ldots,t_{k}]$ $\triangleright$ Knot sequence of $s$
3:  $n\leftarrow\text{length}(T)$
4:  Initialize $dp[0\ldots n-1]$, $prev[0\ldots n-1]$ $\triangleright$ Dynamic programming arrays
5:  $dp[0]\leftarrow 1$, $prev[0]\leftarrow-1$ $\triangleright$ First knot always retained
6:  for $i\leftarrow 1$ to n1n-1 do
7:   $dp[i]\leftarrow\infty$
8:   for $j\leftarrow 0$ to i1i-1 do
9:     if $\text{CheckMergability}(T,j,i,\delta)$ then $\triangleright$ Test if regions between $j$ and $i$ can be merged
10:      $cost\leftarrow dp[j]+1$ $\triangleright$ Retain knot $i$
11:      if $cost<dp[i]$ then
12:        $dp[i]\leftarrow cost$, $prev[i]\leftarrow j$
13:      end if
14:     end if
15:   end for
16:  end for
17:  $new\_knots\leftarrow\text{BacktrackKnots}(prev,T)$
18:  $new\_coeffs\leftarrow\text{ComputeOptimalCoefficients}(T,new\_knots,\delta)$
19:  return $\text{ConstructSpline}(new\_knots,new\_coeffs)$
20:end function
The function $\text{CheckMergability}(T,j,i,\delta)$ verifies whether all regions between knots $j$ and $i$ can be merged into a single polynomial segment with approximation error at most $\delta$. This check can be implemented efficiently using polynomial approximation theory and interval analysis.
5.3 Theoretical Analysis of PolyKAN
We now establish the formal guarantees of the PolyKAN algorithm.
Theorem 3 (Correctness).
For any KAN network $\mathcal{N}$ and error tolerance $\epsilon>0$, the PolyKAN algorithm returns a compressed network $\mathcal{N}^{\prime}$ satisfying $\|\mathcal{N}-\mathcal{N}^{\prime}\|_{\infty}\leq\epsilon$.
Proof.
By construction, each spline compression in Algorithm 2 ensures that the local error is within the allocated budget $\delta_{l}$. The error allocation in Algorithm 1 ensures that the cumulative effect of these local errors satisfies the global bound $\epsilon$, following from the error propagation analysis in Lemma 4. ∎
Theorem 4 (Optimality).
For each individual spline function, Algorithm 2 finds the minimal number of knots that achieves the specified error bound $\delta$.
Proof.
The dynamic programming formulation in Algorithm 2 systematically explores all possible knot sequences. The cost $dp[i]$ represents the minimum number of knots needed to compress the spline up to knot $i$ while satisfying the error bound. The recurrence relation:
$dp[i]=\min_{j<i}\left\{dp[j]+1:\text{regions between $j$ and $i$ are $\delta$-mergable}\right\}$
ensures that all possible merging patterns are considered, guaranteeing optimality for each spline. ∎
Theorem 5 (Complexity).
For a KAN network with $L$ layers, at most $n$ nodes per layer, and at most $k$ knots per spline, the PolyKAN algorithm has time complexity $O(Ln^{2}k^{3})$.
Proof.
Algorithm 2 processes each spline in $O(k^{3})$ time: the outer loop runs $k$ times, the inner loop runs $k$ times, and each mergability check can be performed in $O(k)$ time using appropriate polynomial approximation techniques. With $O(Ln^{2})$ splines in total, the overall complexity is $O(Ln^{2}k^{3})$. ∎
These theoretical guarantees distinguish PolyKAN from heuristic compression methods, providing both performance bounds and formal correctness guarantees.
6 Theoretical Analysis and Case Study
6.1 Computational Complexity Analysis
The $O(Ln^{2}k^{3})$ complexity of PolyKAN is polynomial in all network parameters, making it computationally feasible for practical KAN architectures. The cubic dependence on $k$ is manageable since the number of knots per spline is typically small (10-100 in practice). Furthermore, the algorithm exhibits substantial parallelism, as different splines within and across layers can be compressed independently.
6.2 Error Bound Tightness
A crucial aspect of our theoretical analysis is establishing the tightness of our error bounds.
Proposition 1 (Error Bound Tightness).
There exist KAN networks and input distributions for which the error propagation bounds in Lemma 4 are tight. That is, for any $\delta>0$, there exists a KAN layer such that:
$\|\Phi-\tilde{\Phi}\|_{\infty}=n\cdot\delta$
when $\|s_{j,i}-\tilde{s}_{j,i}\|_{\infty}=\delta$ for all $i,j$.
Proof.
Consider a single KAN layer with $s_{j,i}(x)=c_{j,i}$ (constant functions) and inputs that simultaneously maximize all approximation errors. In this case, the errors accumulate linearly across dimensions, achieving the worst-case bound. ∎
This proposition demonstrates that our error analysis is not overly conservative and provides meaningful guarantees for practical compression scenarios.
6.3 Case Study: Illustrative Example
To elucidate the PolyKAN framework, we present a minimal case study compressing a univariate KAN approximating $f(x)=\sin(2\pi x)$ on $[0,1]$.
Original KAN: Single spline with knots at $[0,0.2,0.4,0.6,0.8,1.0]$ (5 regions).
Compression: With $\epsilon=0.1$, PolyKAN executes as follows:
  1. 1.
    Check mergability of adjacent regions: $0,0.2]\cup[0.2,0.4]$ is $\epsilon$-mergable.
  2. 2.
    Update knot sequence to $[0,0.4,0.6,0.8,1.0]$ (4 regions).
  3. 3.
    Check further merging: $0.4,0.6]\cup[0.6,0.8]$ is $\epsilon$-mergable.
  4. 4.
    Final knot sequence: $[0,0.4,0.8,1.0]$ (3 regions).
This example demonstrates how PolyKAN systematically reduces complexity while maintaining the error guarantee. The dynamic programming approach ensures that this is indeed the optimal compression for the given $\epsilon$.
7 Discussion and Conclusion
7.1 Theoretical Contributions
The PolyKAN framework represents a significant advance in the theory of neural network compression by providing:
  • Formal Guarantees: Unlike heuristic methods, PolyKAN provides provable bounds on both compression ratio and approximation error.
  • Optimality Certificates: The dynamic programming algorithm guarantees minimal compression for each spline function under the given error tolerance.
  • Computational Efficiency: Polynomial-time complexity makes the approach practical for realistic network sizes.
Our work bridges spline theory, polyhedral geometry, and algorithm design, establishing a new paradigm for trustworthy compression of interpretable neural architectures.
7.2 Limitations and Future Directions
While PolyKAN provides strong theoretical foundations, several limitations suggest fruitful directions for future research:
  • Layer-Wise Optimality: The current approach ensures optimal compression for each spline independently, but global optimality across all splines simultaneously remains an open challenge.
  • Error Allocation: Our error propagation bounds, while tight in worst-case, may be conservative for typical inputs. Input-aware error allocation could improve compression ratios.
  • Extended Spline Families: The theory currently focuses on B-splines; extension to other spline families and activation functions would broaden applicability.
Future work will explore these directions, with the goal of developing even more powerful compression frameworks with strong theoretical guarantees.
7.3 Conclusion
We have presented PolyKAN, a comprehensive theoretical framework for KAN compression with provable guarantees. By leveraging the intrinsic polyhedral structure of KANs, we have developed a complete theory of $\epsilon$-equivalent compression, designed an optimal algorithm with polynomial complexity, and established formal guarantees for correctness and minimality. This work provides a foundation for trustworthy compression of interpretable neural architectures, with potential applications in scientific computing, resource-constrained deployment, and theoretical computer science.

### v2

@@
-PolyKAN: A Polyhedral Analysis Framework for Provable and Minimal KAN Compression
+PolyKAN: A Polyhedral Analysis Framework for Provable and Approximately Optimal KAN Compression
 Abstract
-Kolmogorov-Arnold Networks (KANs) have emerged as a promising alternative to traditional Multi-Layer Perceptrons (MLPs), offering enhanced interpretability and a strong mathematical foundation.
+Kolmogorov-Arnold Networks (KANs) have emerged as a promising alternative to traditional Multi-Layer Perceptrons (MLPs), offering enhanced interpretability and a solid mathematical foundation.
 However, their parameter efficiency remains a significant challenge for practical deployment.
 This paper introduces PolyKAN, a novel theoretical framework for KAN compression that provides formal guarantees on both model size reduction and approximation error.
-By leveraging the inherent piecewise polynomial structure of KANs, we formulate the compression problem as one of optimal polyhedral region merging.
-We establish a rigorous polyhedral characterization of KANs, develop a complete theory of $\epsilon$-equivalent compression, and design an optimal dynamic programming algorithm that guarantees minimal compression under specified error bounds.
-Our theoretical analysis demonstrates that PolyKAN achieves provably minimal compression while maintaining strict error control, with polynomial-time complexity in all network parameters.
-The framework provides the first formal foundation for KAN compression with mathematical guarantees, opening new directions for efficient deployment of interpretable neural architectures.
-Keywords: Kolmogorov-Arnold Networks, Model Compression, Polyhedral Analysis, Provable Optimization, Approximation Theory, Computational Complexity
+By leveraging the inherent piecewise polynomial structure of KANs, we formulate the compression problem as a polyhedral region merging task.
+We establish a rigorous polyhedral characterization of KANs, develop a complete theory of $\epsilon$-equivalent compression, and design a dynamic programming algorithm that achieves approximately optimal compression under specified error bounds.
+Our theoretical analysis demonstrates that PolyKAN achieves provably near-optimal compression while maintaining strict error control, with guaranteed global optimality for univariate spline functions.
+This framework provides the first formal foundation for KAN compression with mathematical guarantees, opening new directions for the efficient deployment of interpretable neural architectures.
+Keywords: Kolmogorov-Arnold Networks, Model Compression, Polyhedral Analysis, Provable Optimization, Approximation Theory, Dynamic Programming
 1 Introduction
-The pursuit of interpretable yet powerful neural network architectures has led to renewed interest in mathematically grounded alternatives to traditional black-box models.
+The pursuit of neural network architectures that combine interpretability with strong performance has renewed interest in mathematically grounded alternatives to traditional black-box models.
 Kolmogorov-Arnold Networks (KANs) [1], inspired by the celebrated Kolmogorov-Arnold Representation Theorem [2], have recently emerged as a compelling framework that replaces fixed activation functions with learnable spline transformations.
-While KANs demonstrate remarkable interpretability and empirical performance, their parameter efficiency presents a substantial obstacle to practical adoption, as each network connection requires an independent spline function with multiple parameters.
-Existing approaches to neural network compression—including pruning [3], knowledge distillation [4], and architecture search—largely rely on heuristic strategies without formal guarantees.
-For KANs specifically, the absence of theoretically grounded compression methods represents a significant gap in the literature.
-The spline-based foundation of KANs, however, provides a unique opportunity for rigorous analysis through polyhedral theory, as each spline naturally partitions its input domain into piecewise polynomial regions.
-In this paper, we present PolyKAN, a comprehensive theoretical framework for KAN compression with provable guarantees. Our work is distinguished by three fundamental contributions:
-  1. 1.
-    We establish a complete polyhedral characterization of KANs, demonstrating that their input space partitions exhibit a highly regular axis-aligned structure—in stark contrast to the arbitrary polyhedral complexes of ReLU networks.
+Although KANs demonstrate remarkable interpretability and empirical performance in certain function approximation tasks, their parameter efficiency presents a major obstacle to practical application, as each network connection requires an independent spline function with multiple parameters.
+Existing neural network compression methods—including pruning [3], knowledge distillation [4], and architecture search—largely rely on heuristic strategies and lack formal guarantees.
+For KANs in particular, the absence of theoretically grounded compression methods represents a significant gap in the literature.
+However, the spline-based roots of KANs offer unique opportunities for rigorous analysis through polyhedral theory, as each spline naturally partitions its input domain into piecewise polynomial regions.
+This paper presents PolyKAN, a comprehensive theoretical framework for KAN compression with provable guarantees.
+Our work makes three fundamental contributions: First, we establish a complete polyhedral characterization of KANs, demonstrating that their input space partitions exhibit highly regular axis-aligned structures—in stark contrast to the arbitrary polyhedral complexes of ReLU networks.
 This structural regularity enables rigorous theoretical analysis.
-  2. 2.
-    We develop a formal theory of $\epsilon$-equivalent compression for KANs, providing necessary and sufficient conditions for region merging while maintaining strict error bounds.
-Our analysis includes precise error propagation bounds through multiple network layers.
-  3. 3.
-    We design and analyze an optimal dynamic programming algorithm that guarantees minimal compression under specified error constraints.
-We prove polynomial-time complexity and establish formal guarantees for both correctness and optimality.
-The PolyKAN framework transforms KAN compression from a heuristic process to a mathematically rigorous optimization problem with verifiable guarantees.
-By bridging spline theory, polyhedral geometry, and algorithm design, we provide a foundation for trustworthy compression of interpretable neural architectures.
+Second, we develop a formal theory of $\epsilon$-equivalent compression for KANs, providing necessary and sufficient conditions for region merging while preserving strict error bounds.
+Our analysis includes precise bounds on error propagation through multiple network layers.
+Third, we design and analyze a dynamic programming algorithm that guarantees optimal compression for univariate spline functions and delivers approximately optimal guarantees for the entire network, with provable polynomial time complexity.
+The PolyKAN framework transforms KAN compression from a heuristic process into a mathematically rigorous optimization problem with verifiable guarantees.
+By connecting spline theory, polyhedral geometry, and algorithm design, we lay the foundation for trustworthy compression of interpretable neural architectures.
 2 Background and Related Work
 2.1 Kolmogorov-Arnold Networks and Spline Theory
-The Kolmogorov-Arnold representation theorem [2] states that any multivariate continuous function can be represented as a finite composition of univariate functions and additions.
-KANs [1] operationalize this theorem by placing learnable univariate functions (typically B-splines) on network edges while limiting nodes to summation operations.
-This architectural shift from traditional MLPs provides both mathematical interpretability and empirical advantages in certain function approximation tasks.
+The Kolmogorov-Arnold Representation Theorem [2] states that any multivariate continuous function can be represented as a composition of a finite number of univariate functions and additions.
+KANs [1] implement this theorem by placing learnable univariate functions (typically B-splines) on the network edges while restricting nodes to summation operations.
+This architectural shift from traditional MLPs provides advantages in both mathematical interpretability and empirical performance in certain function approximation tasks.
 Formally, a KAN layer transforms its input $\mathbf{x}\in\mathbb{R}^{n}$ to output $\mathbf{y}\in\mathbb{R}^{m}$ through the relation:
 $y_{j}=\sum_{i=1}^{n}s_{j,i}(x_{i}),\quad j=1,\ldots,m$ (1)
-where each $s_{j,i}:\mathbb{R}\to\mathbb{R}$ is a spline function parameterized by knot positions and polynomial coefficients.
-The approximation properties of splines are well-understood [5], with deep connections to function space theory.
-Recent work [1] has begun exploring the implications of these properties for neural network design, but a comprehensive theoretical framework for KAN optimization remains underdeveloped.
+where each $s_{j,i}:\mathbb{R}\to\mathbb{R}$ is a spline function parameterized by knot locations and polynomial coefficients.
+The approximation properties of splines are well-understood [5] and have deep connections to function space theory.
+Recent work [1] has begun to explore the implications of these properties for neural network design, but a comprehensive theoretical framework for KAN optimization remains underdeveloped.
 2.2 Neural Network Compression
@@
 Neural architecture search [11] automates the design of efficient network structures.
-Despite empirical successes, these approaches largely lack formal guarantees.
-Theoretical work on network compression has typically focused on simplified settings [12] or provided only asymptotic guarantees.
-The lottery ticket hypothesis [10] offers intriguing insights but falls short of providing constructive compression algorithms with bounded error.
+Despite empirical success, most of these methods lack formal guarantees.
+Theoretical work on network compression typically focuses on simplified settings [12] or provides only asymptotic guarantees.
+The lottery ticket hypothesis [10] offers intriguing insights but fails to provide constructive compression algorithms with bounded error.
 2.3 Polyhedral Theory in Deep Learning
 Polyhedral theory provides powerful tools for analyzing piecewise linear neural networks.
-A substantial body of work has studied the linear regions of ReLU networks [6, 7, 8], establishing connections between region counts and network expressivity.
+Substantial work has studied the linear regions of ReLU networks [6, 7, 8], establishing connections between region counts and network expressive power.
 Mixed-integer programming formulations [9] have been developed for verifying properties of ReLU networks.
-However, the polyhedral structure of KANs differs fundamentally from that of ReLU networks.
-While ReLU networks partition their input space with arbitrarily oriented hyperplanes, KAN partitions are axis-aligned due to their spline-based construction.
+Recent work has begun to explore connections between KANs and ReLU networks.
+[13] established formal relationships between piecewise linear KANs and ReLU networks, proving that under certain conditions, both architectures exhibit similar polyhedral complexity.
+Meanwhile, [14] introduced ReLU-KANs, a variant that replaces spline functions with ReLU activations while maintaining the Kolmogorov-Arnold structure, further bridging the gap between these architectures.
+However, the polyhedral structure of general KANs differs fundamentally from that of ReLU networks.
+While ReLU networks partition their input space with hyperplanes of arbitrary orientation, KAN partitions are axis-aligned due to their spline-based construction.
 This structural regularity makes KANs particularly amenable to polyhedral analysis and enables the development of compression algorithms with strong theoretical guarantees.
@@
 3.1 Univariate Spline Partitions
-We begin by formalizing the polyhedral structure of individual spline functions, which form the building blocks of KANs.
+We begin by formalizing the polyhedral structure of individual spline functions, which constitute the building blocks of KANs.
 Definition 1 (Spline Polyhedron).
-For a B-spline function $s:[a,b]\to\mathbb{R}$ with knot sequence $t_{0},t_{1},\ldots,t_{k}$ where $a=t_{0}<t_{1}<\cdots<t_{k}=b$, the spline polyhedron is a tuple $\mathcal{P}_{s}=(R,P)$, where:
-  • $R=\{R_{1},R_{2},\ldots,R_{k}\}$ is a set of regions with $R_{i}=\{x\in\mathbb{R}:t_{i-1}\leq x\leq t_{i}\}$
-  • $P=\{p_{1},p_{2},\ldots,p_{k}\}$ is a set of polynomial functions, with $p_{i}$ defined on $R_{i}$
+For a B-spline function $s:[a,b]\to\mathbb{R}$ defined on interval $[a,b]$ with knot sequence $t_{0},t_{1},\ldots,t_{k}$ where $a=t_{0}<t_{1}<\cdots<t_{k}=b$, its spline polyhedron is a tuple $\mathcal{P}_{s}=(R,P)$ where:
+  • $R=\{R_{1},R_{2},\ldots,R_{k}\}$ is a collection of regions with $R_{i}=\{x\in\mathbb{R}:t_{i-1}\leq x\leq t_{i}\}$
+  • $P=\{p_{1},p_{2},\ldots,p_{k}\}$ is a collection of polynomial functions with $p_{i}$ defined on $R_{i}$
 This definition captures the piecewise polynomial nature of spline functions, where each region corresponds to an interval between consecutive knots.
-3.2 Multi-Layer KAN Partitions
-For multi-layer KANs, the polyhedral structure emerges from the composition of spline functions across layers.
+3.2 Multilayer KAN Partitions
+For multilayer KANs, the polyhedral structure arises from the composition of spline functions across layers.
 Consider an $L$-layer KAN with architecture $n_{0},n_{1},\ldots,n_{L}]$, where $n_{l}$ denotes the number of nodes in layer $l$.
 Theorem 1 (KAN Polyhedral Region Structure).
-The input space partition of an $L$-layer KAN is the refinement of partitions induced by spline functions across all layers. The total number of linear regions satisfies:
+The input space partition of an $L$-layer KAN is a refinement of the partitions induced by spline functions across all layers. The total number of linear regions satisfies:
 $N_{\text{regions}}\leq\prod_{l=1}^{L}\prod_{i=1}^{n_{l-1}}\prod_{j=1}^{n_{l}}(k^{(l)}_{j,i}-1)$ (2)
-where $k^{(l)}_{j,i}$ is the number of knots in spline function $s^{(l)}_{j,i}$ connecting node $i$ in layer l1l-1 to node $j$ in layer $l$.
+where $k^{(l)}_{j,i}$ is the number of knots in the spline function $s^{(l)}_{j,i}$ connecting node $i$ in layer l1l-1 to node $j$ in layer $l$.
 Proof.
 Each spline function $s^{(l)}_{j,i}$ partitions its one-dimensional input space into at most $(k^{(l)}_{j,i}-1)$ intervals.
-Since these partitions act independently along different dimensions at each layer, the Cartesian product of these partitions yields the region structure for that layer’s output space.
-The composition across layers refines these partitions, leading to the product bound.
+Since these partitions act independently along different dimensions at each layer, the Cartesian product of these partitions produces the region structure of that layer’s output space.
+Composition across layers refines these partitions, leading to the product bound.
 ∎
@@
 The polyhedral structure of KANs exhibits three fundamental properties that distinguish them from ReLU networks and enable our compression theory.
-Lemma 1 (Axis-Alignment).
+Lemma 1 (Axis-Alignedness).
 All boundary hyperplanes of KAN polyhedral regions are axis-aligned. That is, each boundary can be expressed as $x_{d}=c$ for some coordinate $d$ and constant $c$.
 Proof.
-The knot points of each spline function $s^{(l)}_{j,i}(x_{i})$ define partitions of the form $x_{i}=t_{m}$, which are hyperplanes orthogonal to the $i$-th coordinate axis.
-The composition of such axis-aligned partitions preserves axis-alignment.
+The knots of each spline function $s^{(l)}_{j,i}(x_{i})$ define partitions of the form $x_{i}=t_{m}$, which are hyperplanes perpendicular to the $i$-th coordinate axis.
+The composition of such axis-aligned partitions preserves axis-alignedness.
 ∎
 Lemma 2 (Rectangular Structure).
-Every KAN polyhedral region is a axis-aligned rectangle (the Cartesian product of intervals).
+Each KAN polyhedral region is an axis-aligned rectangle (Cartesian product of intervals).
 Proof.
-Following from the axis-alignment property, the intersection of half-spaces defined by axis-aligned hyperplanes necessarily produces rectangular regions. ∎
+By the axis-aligned property, intersections of half-spaces defined by axis-aligned hyperplanes necessarily produce rectangular regions. ∎
 Lemma 3 (Function Regularity).
 Within each polyhedral region, a KAN is a smooth multivariate polynomial function.
-Across region boundaries, the KAN maintains continuity (for linear B-splines) or higher-order smoothness (for higher-degree B-splines).
-These structural properties significantly simplify polyhedral analysis compared to ReLU networks, where regions can be arbitrary convex polyhedra and boundaries have arbitrary orientations.
-4 Provable KAN Compression Theory
+Across region boundaries, KANs maintain continuity (for linear B-splines) or higher-order smoothness (for higher-degree B-splines).
+These structural properties significantly simplify polyhedral analysis compared to ReLU networks, whose regions can be arbitrary convex polyhedra with boundaries of arbitrary orientation.
+4 Theory of Provable KAN Compression
 4.1 Formal Problem Statement
-We now formalize the KAN compression problem with provable guarantees.
+We now formalize the problem of KAN compression with provable guarantees.
 Definition 2 ($\epsilon$-Equivalent Compression).
-Given a KAN network $\mathcal{N}:\mathcal{X}\to\mathbb{R}$ and error tolerance $\epsilon>0$, a compressed network $\mathcal{N}^{\prime}$ is an $\epsilon$-equivalent compression of $\mathcal{N}$ if:
+Given a KAN network $\mathcal{N}:\mathcal{X}\to\mathbb{R}$ and an error tolerance $\epsilon>0$, a compressed network $\mathcal{N}^{\prime}$ is an $\epsilon$-equivalent compression of $\mathcal{N}$ if:
 $\|\mathcal{N}-\mathcal{N}^{\prime}\|_{\infty}=\sup_{x\in\mathcal{X}}|\mathcal{N}(x)-\mathcal{N}^{\prime}(x)|\leq\epsilon$ (3)
-Our objective is to find an $\epsilon$-equivalent compression $\mathcal{N}^{\prime}$ with minimal number of parameters (or equivalently, minimal total knots across all spline functions).
-4.2 Region Mergability Theory
-The foundation of our compression approach is the theory of merging adjacent polyhedral regions while controlling approximation error.
+Problem 1 (Optimal KAN Compression).
+Given a KAN network $\mathcal{N}$ and an error tolerance $\epsilon>0$, find an $\epsilon$-equivalent compression $\mathcal{N}^{\prime}$ that minimizes the total number of knots across all spline functions.
+4.2 Computational Complexity Analysis
+We first analyze the computational complexity of the optimal KAN compression problem.
+Theorem 2 (NP-Hardness).
+The optimal KAN compression problem is NP-hard.
+Proof.
+Consider a simplified version: given a set of one-dimensional intervals $I_{1},I_{2},\ldots,I_{n}$ with corresponding polynomials $p_{1},p_{2},\ldots,p_{n}$, and an error tolerance $\epsilon$, find the smallest set of knots such that each merged interval can be approximated by a single polynomial with error $\leq\epsilon$.
+This can be reduced to the set cover problem, which is classically NP-complete.
+More specifically, given a set cover instance $(U,\mathcal{S})$ where $U$ is the universe and $\mathcal{S}$ is a family of subsets, we can construct a KAN compression instance by mapping each element in $U$ to an interval and each subset in $\mathcal{S}$ to a feasible interval merging.
+The optimal compression corresponds to the minimum set cover.
+∎
+Although the overall problem is NP-hard, the special structure of KANs allows us to design efficient approximation algorithms.
+4.3 Theory of Region Mergability
+The foundation of our compression approach is a theory of merging adjacent polyhedral regions while controlling approximation error.
 Definition 3 (Region Mergability).
-Let $R_{i}$ and $R_{j}$ be two adjacent polyhedral regions in a KAN, with corresponding polynomial functions $p_{i}$ and $p_{j}$.
-The regions are $\epsilon$-mergable if there exists a single polynomial $p_{ij}$ such that:
+Let $R_{i}$ and $R_{j}$ be two adjacent polyhedral regions in a KAN with corresponding polynomial functions $p_{i}$ and $p_{j}$.
+These regions are $\epsilon$-mergable if there exists a single polynomial $p_{ij}$ such that:
 $\max\left\{\max_{x\in R_{i}}|p_{i}(x)-p_{ij}(x)|,\max_{x\in R_{j}}|p_{j}(x)-p_{ij}(x)|\right\}\leq\epsilon$ (4)
-This definition captures the intuition that we can replace two different polynomials defined on adjacent regions with a single polynomial that approximates both within tolerance $\epsilon$.
+This definition captures the intuition that we can replace two different polynomials defined on adjacent regions with a single polynomial that approximates both original polynomials within tolerance $\epsilon$.
 Theorem 3 (Knot Elimination Condition).
-Let $t_{m}$ be an interior knot of spline function $s(x)$, with adjacent regions $R_{m-1}=[t_{m-1},t_{m}]$ and $R_{m}=[t_{m},t_{m+1}]$.
-If $R_{m-1}$ and $R_{m}$ are $\epsilon$-mergable, then knot $t_{m}$ can be eliminated while maintaining $\epsilon$-equivalence.
+Let $t_{m}$ be an interior knot of a spline function $s(x)$ with adjacent regions $R_{m-1}=[t_{m-1},t_{m}]$ and $R_{m}=[t_{m},t_{m+1}]$.
+If $R_{m-1}$ and $R_{m}$ are $\epsilon$-mergable, then knot $t_{m}$ can be eliminated while preserving $\epsilon$-equivalence.
 Proof.
 If $R_{m-1}$ and $R_{m}$ are $\epsilon$-mergable, there exists a polynomial $p$ such that $|s(x)-p(x)|\leq\epsilon$ for all $x\in R_{m-1}\cup R_{m}$.
-By replacing the original piecewise representation with $p$ on the merged region $t_{m-1},t_{m+1}]$, we obtain a new spline function $\tilde{s}$ that satisfies $\|s-\tilde{s}\|_{\infty}\leq\epsilon$ while eliminating one knot.
+By replacing the original piecewise representation on the merged region $t_{m-1},t_{m+1}]$ with $p$, we obtain a new spline function $\tilde{s}$ satisfying $\|s-\tilde{s}\|_{\infty}\leq\epsilon$ while eliminating one knot.
 ∎
 4.4 Error Propagation Analysis
-For multi-layer KAN compression, we must understand how compression errors propagate through the network.
+For multilayer KAN compression, we must understand how compression errors propagate through the network.
 Lemma 4 (Single-Layer Error Propagation).
-Consider a KAN layer $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$ with:
+Consider a KAN layer $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$ where:
 $y_{j}=\sum_{i=1}^{n}s_{j,i}(x_{i}),\quad j=1,\ldots,m$
@@
 $\leq\sum_{i=1}^{n}\delta=n\delta$
-Taking the supremum over all $\mathbf{x}$ and maximum over $j$ completes the proof. ∎
-This lemma enables hierarchical error budgeting across network layers.
+Taking the supremum over all $\mathbf{x}$ and the maximum over $j$ completes the proof. ∎
+This lemma enables layered error budget allocation across network layers.
 Given a global error tolerance $\epsilon$, we can allocate error budgets $\delta_{l}$ to each layer $l$ such that the cumulative effect satisfies the global bound.
-4.4 Optimal Compression Formulation
-We now formulate KAN compression as a combinatorial optimization problem.
-Let $z_{l,i,j,m}\in\{0,1\}$ indicate whether the $m$-th knot of spline $s^{(l)}_{j,i}$ is retained, and let $\delta_{l,i,j}$ be the error budget allocated to this spline.
-The optimal compression problem becomes:
-$\min$ $\sum_{l=1}^{L}\sum_{i=1}^{n_{l-1}}\sum_{j=1}^{n_{l}}\sum_{m=1}^{k_{l,i,j}}z_{l,i,j,m}$ (5)
-s.t. $\epsilon$-equivalence constraints
-Knot sequence feasibility
-$z_{l,i,j,m}\in\{0,1\}$
-While this integer programming formulation captures the optimality structure, its direct solution is computationally challenging.
-In the next section, we develop an efficient dynamic programming algorithm that exploits the problem structure to achieve optimality with polynomial complexity.
-5 The PolyKAN Compression Algorithm
-5.1 Algorithm Overview
-We present the PolyKAN compression algorithm, which systematically reduces KAN complexity while guaranteeing $\epsilon$-equivalence.
-The algorithm operates layer by layer, compressing each spline function optimally using dynamic programming.
-Algorithm 1 PolyKAN Provable Compression Algorithm
-1:Original KAN network $\mathcal{N}$, global error bound $\epsilon$
-2:Compressed KAN network $\mathcal{N}^{\prime}$ with $\|\mathcal{N}-\mathcal{N}^{\prime}\|_{\infty}\leq\epsilon$
-[…]
-4:for $l\leftarrow 1$ to $L$ do
-5:  for each spline function $s$ in layer $l$ do
-6:   $\tilde{s}\leftarrow\text{OptimalSplineCompression}(s,\delta_{l})$ $\triangleright$ Algorithm 2
+5 Approximately Optimal Compression Algorithms
+5.1 Optimal Compression for Univariate Splines
+Although overall KAN compression is NP-hard, for univariate spline functions we can design efficient algorithms with optimality guarantees.
+Algorithm 1 Optimal Compression for Single Spline
+1:Spline function $s$, knot sequence $t_{0},t_{1},\ldots,t_{k}$, error tolerance $\epsilon$
+2:Compressed knot sequence
+3:Initialize DP table: $dp[i]\leftarrow i+1$ for $i=0,1,\ldots,k$ $\triangleright$ $dp[i]$: min knots to reach $t_{i}$
+4:Initialize backtrack pointers: $prev[i]\leftarrow-1$ for $i=0,1,\ldots,k$
+5:for $i=1$ to $k$ do
+6:  for $j=0$ to i1i-1 do
+7:   Check if interval $t_{j},t_{i}]$ can be approximated by a single polynomial $p$ with error $\leq\epsilon$
+8:   if mergable and $dp[j]+1<dp[i]$ then
+9:     $dp[i]\leftarrow dp[j]+1$
+10:     $prev[i]\leftarrow j$
+11:   end if
+12:  end for
+13:end for
+14:return Optimal knot sequence constructed by backtracking prevprev array
+Theorem 4 (Univariate Optimality).
+Algorithm 1 guarantees the global optimal solution for the single spline optimal compression problem.
+Proof.
+The algorithm is a classic interval partitioning dynamic programming.
+The optimal substructure property holds: the optimal solution from $t_{0}$ to $t_{i}$ must consist of the optimal solution from $t_{0}$ to some $t_{j}$ plus a single polynomial approximation for interval $t_{j},t_{i}]$.
+The dynamic programming correctly explores all possible partition points.
+∎
+Theorem 5 (Time Complexity).
+Algorithm 1 has time complexity $O(k^{3}\cdot T_{\text{fit}})$, where $k$ is the original number of knots and $T_{\text{fit}}$ is the time for polynomial fitting.
+Proof.
+The algorithm has $O(k^{2})$ state transitions, each requiring $O(k)$ time to check the feasibility of interval merging (via polynomial fitting and error computation), yielding total complexity $O(k^{3}\cdot T_{\text{fit}})$.
+∎
+5.2 Approximate Compression for Multilayer KANs
+For entire KAN networks, we employ a layered compression strategy.
+Algorithm 2 Approximate Compression for Multilayer KAN
+1:KAN network $\mathcal{N}$, global error tolerance $\epsilon$
+2:Compressed KAN network $\mathcal{N}^{\prime}$
+3:Allocate global error budget $\epsilon$ proportionally across layers: $\epsilon=\sum_{l=1}^{L}\epsilon_{l}$
+4:for each layer $l=1$ to $L$ do
+5:  for each spline function $s$ in this layer do
+6:   Compress spline $s$ using Algorithm 1 with error budget $\epsilon_{l}/n_{l}$
 7:  end for
 8:end for
-[…]
-10:return $\mathcal{N}^{\prime}$
-The error allocation in Step 1 distributes the global error budget $\epsilon$ across layers according to the error propagation analysis, ensuring that the cumulative effect satisfies the global bound.
-5.2 Optimal Spline Compression
-The core of PolyKAN is the optimal compression of individual spline functions via dynamic programming.
-Algorithm 2 OptimalSplineCompression
-1:function OptimalSplineCompression($s$, $\delta$)
-2:  $T\leftarrow[t_{0},t_{1},\ldots,t_{k}]$ $\triangleright$ Knot sequence of $s$
-3:  $n\leftarrow\text{length}(T)$
-4:  Initialize $dp[0\ldots n-1]$, $prev[0\ldots n-1]$ $\triangleright$ Dynamic programming arrays
-5:  $dp[0]\leftarrow 1$, $prev[0]\leftarrow-1$ $\triangleright$ First knot always retained
-6:  for $i\leftarrow 1$ to n1n-1 do
-7:   $dp[i]\leftarrow\infty$
-8:   for $j\leftarrow 0$ to i1i-1 do
-9:     if $\text{CheckMergability}(T,j,i,\delta)$ then $\triangleright$ Test if regions between $j$ and $i$ can be merged
-10:      $cost\leftarrow dp[j]+1$ $\triangleright$ Retain knot $i$
-11:      if $cost<dp[i]$ then
-12:        $dp[i]\leftarrow cost$, $prev[i]\leftarrow j$
-13:      end if
-14:     end if
-15:   end for
-16:  end for
-17:  $new\_knots\leftarrow\text{BacktrackKnots}(prev,T)$
-18:  $new\_coeffs\leftarrow\text{ComputeOptimalCoefficients}(T,new\_knots,\delta)$
-19:  return $\text{ConstructSpline}(new\_knots,new\_coeffs)$
-20:end function
-The function $\text{CheckMergability}(T,j,i,\delta)$ verifies whether all regions between knots $j$ and $i$ can be merged into a single polynomial segment with approximation error at most $\delta$.
-This check can be implemented efficiently using polynomial approximation theory and interval analysis.
-5.3 Theoretical Analysis of PolyKAN
-We now establish the formal guarantees of the PolyKAN algorithm.
-Theorem 3 (Correctness).
-For any KAN network $\mathcal{N}$ and error tolerance $\epsilon>0$, the PolyKAN algorithm returns a compressed network $\mathcal{N}^{\prime}$ satisfying $\|\mathcal{N}-\mathcal{N}^{\prime}\|_{\infty}\leq\epsilon$.
+9:return compressed network $\mathcal{N}^{\prime}$
+Theorem 6 (Approximation Guarantee).
+The compressed network $\mathcal{N}^{\prime}$ produced by Algorithm 2 satisfies $\|\mathcal{N}-\mathcal{N}^{\prime}\|_{\infty}\leq\epsilon$.
 Proof.
-By construction, each spline compression in Algorithm 2 ensures that the local error is within the allocated budget $\delta_{l}$.
-The error allocation in Algorithm 1 ensures that the cumulative effect of these local errors satisfies the global bound $\epsilon$, following from the error propagation analysis in Lemma 4.
+By the single-layer error propagation lemma, each layer introduces at most $n_{l}\cdot(\epsilon_{l}/n_{l})=\epsilon_{l}$ error.
+The accumulation of errors across layers is controlled through error budget allocation, ensuring total error does not exceed $\sum_{l=1}^{L}\epsilon_{l}=\epsilon$.
 ∎
-Theorem 4 (Optimality).
-For each individual spline function, Algorithm 2 finds the minimal number of knots that achieves the specified error bound $\delta$.
+5.3 Optimality Gap Analysis
+Although layered compression cannot guarantee global optimality, we can quantify its optimality gap.
+Theorem 7 (Approximation Ratio).
+Under the assumption of uniform error budget allocation, the gap between the compression ratio achieved by Algorithm 2 and that of the optimal solution is bounded polynomially by the network depth and width.
 Proof.
-The dynamic programming formulation in Algorithm 2 systematically explores all possible knot sequences.
-The cost $dp[i]$ represents the minimum number of knots needed to compress the spline up to knot $i$ while satisfying the error bound.
-The recurrence relation:
-$dp[i]=\min_{j<i}\left\{dp[j]+1:\text{regions between $j$ and $i$ are $\delta$-mergable}\right\}$
-ensures that all possible merging patterns are considered, guaranteeing optimality for each spline. ∎
-Theorem 5 (Complexity).
-For a KAN network with $L$ layers, at most $n$ nodes per layer, and at most $k$ knots per spline, the PolyKAN algorithm has time complexity $O(Ln^{2}k^{3})$.
-Proof.
-Algorithm 2 processes each spline in $O(k^{3})$ time: the outer loop runs $k$ times, the inner loop runs $k$ times, and each mergability check can be performed in $O(k)$ time using appropriate polynomial approximation techniques.
-With $O(Ln^{2})$ splines in total, the overall complexity is $O(Ln^{2}k^{3})$.
+Let $OPT$ be the number of knots in the globally optimal compression and $ALG$ be the number obtained by the algorithm.
+Since each layer is compressed independently optimally and error propagation is linear, the optimality gap is bounded by the optimality of error budget allocation across layers.
+Specifically, there exists a constant $C$ (dependent on network structure) such that $ALG\leq C\cdot OPT$.
 ∎
-These theoretical guarantees distinguish PolyKAN from heuristic compression methods, providing both performance bounds and formal correctness guarantees.
-6 Theoretical Analysis and Case Study
-6.1 Computational Complexity Analysis
-The $O(Ln^{2}k^{3})$ complexity of PolyKAN is polynomial in all network parameters, making it computationally feasible for practical KAN architectures.
-The cubic dependence on $k$ is manageable since the number of knots per spline is typically small (10-100 in practice).
-Furthermore, the algorithm exhibits substantial parallelism, as different splines within and across layers can be compressed independently.
-6.2 Error Bound Tightness
-A crucial aspect of our theoretical analysis is establishing the tightness of our error bounds.
-Proposition 1 (Error Bound Tightness).
-There exist KAN networks and input distributions for which the error propagation bounds in Lemma 4 are tight. That is, for any $\delta>0$, there exists a KAN layer such that:
-$\|\Phi-\tilde{\Phi}\|_{\infty}=n\cdot\delta$
-when $\|s_{j,i}-\tilde{s}_{j,i}\|_{\infty}=\delta$ for all $i,j$.
-Proof.
-Consider a single KAN layer with $s_{j,i}(x)=c_{j,i}$ (constant functions) and inputs that simultaneously maximize all approximation errors.
-In this case, the errors accumulate linearly across dimensions, achieving the worst-case bound.
-∎
-This proposition demonstrates that our error analysis is not overly conservative and provides meaningful guarantees for practical compression scenarios.
-6.3 Case Study: Illustrative Example
-To elucidate the PolyKAN framework, we present a minimal case study compressing a univariate KAN approximating $f(x)=\sin(2\pi x)$ on $[0,1]$.
-Original KAN: Single spline with knots at $[0,0.2,0.4,0.6,0.8,1.0]$ (5 regions).
-Compression: With $\epsilon=0.1$, PolyKAN executes as follows:
-  1. 1.
-    Check mergability of adjacent regions: $0,0.2]\cup[0.2,0.4]$ is $\epsilon$-mergable.
-  2. 2.
-    Update knot sequence to $[0,0.4,0.6,0.8,1.0]$ (4 regions).
-  3. 3.
-    Check further merging: $0.4,0.6]\cup[0.6,0.8]$ is $\epsilon$-mergable.
-  4. 4.
-    Final knot sequence: $[0,0.4,0.8,1.0]$ (3 regions).
-This example demonstrates how PolyKAN systematically reduces complexity while maintaining the error guarantee.
-The dynamic programming approach ensures that this is indeed the optimal compression for the given $\epsilon$.
-7 Discussion and Conclusion
-7.1 Theoretical Contributions
-The PolyKAN framework represents a significant advance in the theory of neural network compression by providing:
-  • Formal Guarantees: Unlike heuristic methods, PolyKAN provides provable bounds on both compression ratio and approximation error.
-  • Optimality Certificates: The dynamic programming algorithm guarantees minimal compression for each spline function under the given error tolerance.
-  • Computational Efficiency: Polynomial-time complexity makes the approach practical for realistic network sizes.
-Our work bridges spline theory, polyhedral geometry, and algorithm design, establishing a new paradigm for trustworthy compression of interpretable neural architectures.
-7.2 Limitations and Future Directions
-While PolyKAN provides strong theoretical foundations, several limitations suggest fruitful directions for future research:
-  • Layer-Wise Optimality: The current approach ensures optimal compression for each spline independently, but global optimality across all splines simultaneously remains an open challenge.
-  • Error Allocation: Our error propagation bounds, while tight in worst-case, may be conservative for typical inputs. Input-aware error allocation could improve compression ratios.
-  • Extended Spline Families: The theory currently focuses on B-splines; extension to other spline families and activation functions would broaden applicability.
-Future work will explore these directions, with the goal of developing even more powerful compression frameworks with strong theoretical guarantees.
-7.3 Conclusion
-We have presented PolyKAN, a comprehensive theoretical framework for KAN compression with provable guarantees.
-By leveraging the intrinsic polyhedral structure of KANs, we have developed a complete theory of $\epsilon$-equivalent compression, designed an optimal algorithm with polynomial complexity, and established formal guarantees for correctness and minimality.
-This work provides a foundation for trustworthy compression of interpretable neural architectures, with potential applications in scientific computing, resource-constrained deployment, and theoretical computer science.
+6 Conclusion and Future Work
+We have presented PolyKAN, a theoretical framework for KAN compression with provable guarantees.
+Although optimal KAN compression is NP-hard, we leveraged the axis-aligned structure of KANs to design efficient dynamic programming algorithms that guarantee optimality for univariate splines and provide approximation guarantees for the entire network.
+Future work includes developing improved algorithms with better approximation ratios, investigating information-theoretic lower bounds for KAN compression, and extending the framework to other types of spline functions and network architectures.